OpenStudy (anonymous):

sqrt(x+4)*sqrt(x+4) = x + 4 right?

OpenStudy (lgbasallote):

\[\LARGE \implies (\sqrt{x + 4}) \times (\sqrt{x+4}) \implies \sqrt{(x+4)\times (x+4)}\] \[\LARGE\implies \sqrt{(x+4)^2} \implies x +4\]

OpenStudy (lgbasallote):

better @Leaper ?

OpenStudy (anonymous):

I don't understand why it got cutt off

OpenStudy (lgbasallote):

do you get the process though?

OpenStudy (anonymous):

Igba Do u have Twiddla

OpenStudy (lgbasallote):

no..

OpenStudy (callisto):

\[\sqrt{(x+4)} \times \sqrt{(x+4)}=(x+4)^{\frac{1}{2}} \times (x+4)^{\frac{1}{2}}\]\[ = (x+4)^{\frac{1}{2}+\frac{1}{2}} = (x+4)^1 = x+4\]

OpenStudy (anonymous):

\[ \sqrt{x + 4} \times \sqrt{x+4} =\sqrt{16+8 x+x^2}=\sqrt{(4+x)^2} = \pm(x+4) \]

OpenStudy (anonymous):

Is sqrt(x^2 + 8x + 16) = x + 4

myininaya (myininaya):

\[\sqrt{x+4} \cdot \sqrt{x+4}=(x+4)=x+4\] but \[\sqrt{(x+4)^2}=\pm (x+4)\]

OpenStudy (lgbasallote):

or \[\huge \sqrt{(x+4)^2} \implies (x+4)^{\frac{2}{2}} = (x+4)^1 \implies x + 4\]

myininaya (myininaya):

ffm i disagree with what yu have

OpenStudy (anonymous):

\[ \sqrt{x + 4} \times \sqrt{x+4} =\sqrt{16+8 x+x^2}\]\[=\sqrt{(4+x)^2} = \pm(x+4) \]

myininaya (myininaya):

\[\sqrt{x+4} \ge 0\] it cannot be negative

myininaya (myininaya):

\[\sqrt{x+4} \cdot \sqrt{x+4} =(x+4)\]

myininaya (myininaya):

but if we have \[\sqrt{(x+4)^2}=\pm (x+4)\]

myininaya (myininaya):

\[\sqrt{x+4} \cdot \sqrt{x+4} \neq \pm (x+4) \]

OpenStudy (anonymous):

I don't understand why sqrt(x^2 + 8x + 16) = x + 4

myininaya (myininaya):

it just equals the positive output

myininaya (myininaya):

Do you know how to factor x^2+8x+16?

OpenStudy (anonymous):

Yes I do

myininaya (myininaya):

what do you get

OpenStudy (callisto):

If it is \( y^2 = (x+4)^2\) , then \(y = \pm (x+4)\) but the sqrt there is already positive, why should we make it negative? If it is \(\pm \sqrt{(x+4)^2}\), then it is \(\pm (x+4)\) Hmm...

OpenStudy (anonymous):

You get (x+4)^2

myininaya (myininaya):

\[\sqrt{(x+4)^2}=|x+4|=\pm(x+4)\]

myininaya (myininaya):

By definition of course :)

OpenStudy (lgbasallote):

why is there a minus @myininaya ? doesnt the square root just cancel out thus canceling the \(\pm\) tOO?

OpenStudy (anonymous):

I just don't understand this concept, I don't know how to view sqrt(x+2)*sqrt(x+2) = x+2

OpenStudy (callisto):

*totally new to me* Then FoolForMath is correct?!

OpenStudy (lgbasallote):

\[\LARGE \sqrt a \times \sqrt a = a \longleftarrow \text{rule}\]

mathslover (mathslover):

\[{\sqrt{x+4}*\sqrt{x+4}=\sqrt{(x+4})^2}\] \[\huge{x+4}\]

OpenStudy (lgbasallote):

just accept that fact for now @Leaper no need to understand the philosophies as to how, why, etc ;)

mathslover (mathslover):

i think \(\pm{x+4}\) is not possible

myininaya (myininaya):

Ok if we have \[\sqrt{a} \sqrt{a}\] Then the answer is a If we have \[\sqrt{a^2}\] Then the answer is \[\sqrt{a^2}=|a|=\pm a\]

mathslover (mathslover):

according to the rule as lgba said

myininaya (myininaya):

of course assuming a>0 :)

myininaya (myininaya):

for that first one i wrote

OpenStudy (anonymous):

What if I have sqrt(10x+50y+20z+30a)*sqrt(10x+50y+20z+30a) does it equal 10x+50y+20z+30a

OpenStudy (lgbasallote):

yep

mathslover (mathslover):

@myininaya will it be \(\sqrt{a}^2 \) instead of\(\sqrt{a^2}\)

OpenStudy (lgbasallote):

isnt \((\sqrt x)^2 = \sqrt{x^2}\) lol

myininaya (myininaya):

which one are you talking about @mathslover

myininaya (myininaya):

no @lgbasallote

mathslover (mathslover):

Ok if we have a√a√ Then the answer is a If we have a2−−√ Then the answer is a2−−√=|a|=±a for this comment

OpenStudy (anonymous):

Igba the rule makes sense when sqrt(3x)*sqrt(3x) = 3x because sqrt(3*3*x*x) = sqrt(9x^2) = 3x

OpenStudy (lgbasallote):

yup..you see now? :)

OpenStudy (anonymous):

I understand when there is one term not multiple terms under a radical

OpenStudy (anonymous):

Whatever I wil accept the rule and try to figure it out later

OpenStudy (lgbasallote):

haha that's right!! i accepted that rule and never complained =))))

myininaya (myininaya):

which one is the problem what rule?

OpenStudy (anonymous):

I'm trying to understand it on a deeper level

myininaya (myininaya):

\[(\sqrt{a})^2=\sqrt{a} \cdot \sqrt{a}=a , a>0\] \[\sqrt{a^2}=|a|=\pm(a)\]

OpenStudy (lgbasallote):

the rule that \[\sqrt a \sqrt a = a\] what astounds us though is this new revelation that \(\sqrt{a^2} = \pm a\)

OpenStudy (anonymous):

I will tell you my perspective sqrt(a) = a^1/2, sqrt(a) = a^1/2 When you multiply exponentsn you add the exponents, 1/2 + 1/2 = 2/2 = 1 exponent a^1/1 = a right?

OpenStudy (anonymous):

but when sqrt(x+5) = x^1/2 + 5^1/2

myininaya (myininaya):

no no you can't do that leaper

OpenStudy (lgbasallote):

sqrt (x+5) is NOT equal to x^1/2 + 5^1/2

OpenStudy (anonymous):

Why

mathslover (mathslover):

\[\sqrt{a+b}=\sqrt{a}+\sqrt{b}\] that is not correct

OpenStudy (lgbasallote):

imagine this \[\sqrt{9+ 16}\] that is originally equal to \(\sqrt{25} = 5\) but if you try your method you come up with \[\large \sqrt{3^2 + 4^2} = \sqrt{3^2} + \sqrt{4^2} = 3 + 4 = 7\] and that is NOT equal to 5

OpenStudy (lgbasallote):

you get why you cant do it?

mathslover (mathslover):

let \(\sqrt{a+b}=x\) square both sides ..\(a+b=x^2\) according to you \(\sqrt{a}+\sqrt{b}=x\) again square both sides ... \(a+b+2\sqrt{ab}=x^2\) that is not equal to a+b=x^2 hence ur statement is proved as wrong

mathslover (mathslover):

@lgbasallote has put constants there in the place of a and b that acted as a good example for you hope you now understood ?

myininaya (myininaya):

@lgbasallote I will say more about that rule: \[\sqrt{x^2}=\pm (x)\] \[\sqrt{x^2}=-x, x<0\] \[\sqrt{x^2}=x , x>0\] \[\sqrt{x^2}=0, x=0\] So Lets test this out I said if x<0 then sqrt(x^2)=-x So what is a number less than 0? How about -4 Does that part of the rule hold? \[\sqrt{(-4)^2}=-(-4)\] yes :) We get a positive number because -(-4)=4 :)

OpenStudy (anonymous):

Igba so all calculations inside a radical have to be done first before square rooting take action?

OpenStudy (lgbasallote):

yup

mathslover (mathslover):

yahooo !!!

mathslover (mathslover):

i am very happy today .... my all LaTeX went crashed..

OpenStudy (anonymous):

I think that's that info I was looking for

OpenStudy (lgbasallote):

oh goodie :DDD i was prepared for more examples but if you get it then it's all yahoo

OpenStudy (anonymous):

Igba so in sqrt(x+4), 4 has to be added to x before the sum of those two gets square rooted right?

OpenStudy (lgbasallote):

yup...assuming you have a value of x of course

OpenStudy (lgbasallote):

if there's no numerical value to x then you cant add it to 4

myininaya (myininaya):

Here is another example showing the statement \[\sqrt{x+y}=\sqrt{x}+\sqrt{y}\] is false: \[\sqrt{2}=\sqrt{1+1} \neq \sqrt{1} + \sqrt{1}=1+1=2\]

OpenStudy (anonymous):

Thank guys.

OpenStudy (anonymous):

Igba regardless if there is a numerical value the x+4 inside sqrt has to be done before square rooting right

OpenStudy (anonymous):

like hypothetically speaking

OpenStudy (lgbasallote):

yepp

mathslover (mathslover):

there is a also a gud question that why is \(\sqrt{-1}=i\) andnot \(\pm i\)

mathslover (mathslover):

i think this is very true was also told by my maths teacher once but when i asked the teacher he went away .. as a function it is always + sqrt(25) is +5 but in an equation like x^2 = 5, you will have two solutions +sqrt(5) and - sqrt(5)

OpenStudy (anonymous):

Okay finally I understand the problem, the correct way to solve this one is \[ \sqrt{x + 4} \times \sqrt{x+4} =\sqrt{16+8 x+x^2} \] \[ =\sqrt{(4+x)}^2 \color{Red}{\neq \sqrt{(4+x)^2} } = (x+4)\]

mathslover (mathslover):

we should clear this doubt to the maths inventor ...

OpenStudy (anonymous):

More importantly \[ \sqrt{x} \times \sqrt{x}=(\sqrt{x})^2\neq\sqrt{(x)^2} \forall x \in \mathbb{R}\] But then the OP didn't defined \(x\), we are all assuming it to be real.

OpenStudy (anonymous):

And yes roting should be done afterwards.